OpenStudy (anonymous):

One corner of a triangle has a 60° angle and the length of the two adjacent sides are in ratio 1 : 3. Calculate the angles of the other triangle corners (0,1°:s precision, 1 point / correct angle)

OpenStudy (anonymous):

use the rule of ratios since one corner is 60 degree you have (180-60=120) to work with therefore for the portion which 1 share in 1:3 1/4 x 120=60 and the other portion with 3 share 3/4 x 120=90